OpenStudy (anonymous):

A curve with a y-intercept of 4 and passing through the point (2,e^2) has the property that the slope of the curve at any point (x,y) is directly proportional to the y-coordinate of the point. What is an equation of this curve?

OpenStudy (ribhu):

Let (dy/dx)= Ky on rearrangement (dy/y)=Kdx now integrating both the sides we get lny=kx+c -------- (1) since the curve passes through (2,e^2) with intercepts, substituting y=e^2, x=2 and intercept 'c'= 4 we get the equation of the curve as x+lny=2 (Answer)